OpenStudy (anonymous):

Solve 2x2 = 14x + 20.

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

or is it x=2 or x=5

OpenStudy (anonymous):

0=10x+20 x = -2

OpenStudy (anonymous):

x=-2 or -5?

OpenStudy (callisto):

2x^2 = 14x + 20 2x^2 - 14x - 20 =0 Use quadratic formula to get the answer. PS: it's not -2

OpenStudy (anonymous):

2x2=14x + 20 is same as 4x = 14x + 20

OpenStudy (callisto):

I think the asker mean \(2x^2=14x + 20\)